OpenStudy (anonymous):

What are the solutions for |x - 4 = 3|?

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

Question should be like this |x - 4| = 3 so \[x - 4 = \pm3\] i.e. x - 4 = 3 or x - 4 = -3 i.e. x = 4+3 or x = 4-3 i.e. x = 7 or x = 1 so x=1, 7

OpenStudy (anonymous):

Ahh. Now I get it. Okay, thank you!!